according to the following how many moles of sulfurous acid (H2SO3) will be formed upon the complete reaction of 30.0 grams of sulfur dioxide with excess water? SO2+H2O(l)H2SO3(g)
Answer #1

mole of SO2 = 30/64 mol

now mole ratio of SO2 to H2SO3 is 1:1

so moles of H2SO3 formed = 30/64 = 0.469 mol

= 0.47 mol